# CleanMel: Mel-Spectrogram Enhancement for Improving Both Speech Quality and ASR

### v1

@@ V Results and Analyses
In this section, we present and analyze the speech enhancement performance and ASR results, and compare the model size and computational complexity as well.
@@ V-A Speech enhancement results
TABLE IV: DNSMOS scores on real test sets.
Mode Enhancement method CHiME4 real EARS REVERB real DNS w. reverb. RealMAN static
P.835 P.808 P.835 P.808 P.835 P.808 P.835 P.808 P.835 P.808
OVL SIG BAK OVL SIG BAK OVL SIG BAK OVL SIG BAK OVL SIG BAK
unproc. 1.37 1.99 1.34 2.43 2.02 2.94 2.04 2.80 1.39 1.74 1.53 2.81 1.42 1.89 1.50 2.74 1.53 2.03 1.75 2.56
Online FullSubNet [15] 2.41 3.05 3.08 3.22 2.73 3.12 3.55 3.52 2.72 3.13 3.60 3.61 2.41 2.98 3.15 3.36 2.49 2.97 3.33 3.12
Demucs [50] 2.84 3.12 3.90 3.26 2.94 3.21 3.94 3.38 2.96 3.25 3.94 3.50 2.50 2.82 3.66 3.17 -
oSpatialNet [22] 2.61 3.15 3.36 3.52 2.85 3.22 3.65 3.65 3.03 3.33 3.93 3.84 2.79 3.16 3.70 3.67 2.79 3.19 3.66 3.46
CleanMel-S-map (ours) 2.95 3.26 3.86 3.77 2.89 3.24 3.69 3.72 3.17 3.45 4.02 3.94 3.08 3.36 3.97 3.84 2.90 3.24 3.81 3.52
CleanMel-S-mask (ours) 2.74 3.26 3.43 3.59 2.83 3.19 3.64 3.70 3.07 3.40 3.91 3.80 2.95 3.30 3.81 3.75 2.73 3.18 3.53 3.36
Offline FullSubNet [15] 2.39 2.93 3.18 3.31 2.47 3.10 3.11 3.27 2.71 3.11 3.62 3.62 2.74 3.16 3.51 3.54 2.47 2.97 3.27 3.13
VoiceFixer [19] 2.95 3.25 3.89 3.65 2.94 3.25 3.83 3.55 2.88 3.26 3.61 3.62 3.11 3.40 3.97 3.84 -
StoRM [51] 3.29 3.57 4.03 3.87 2.97 3.37 3.65 3.79 3.25 3.53 4.01 4.01 3.25 3.53 4.03 3.87 3.04 3.42 3.80 3.68
SpatialNet [21] 2.76 3.32 3.39 3.51 2.86 3.27 3.58 3.59 3.07 3.41 3.87 3.84 2.91 3.33 3.68 3.65 2.88 3.32 3.66 3.45
SpatialNet-Mamba 2.94 3.35 3.71 3.65 2.94 3.31 3.69 3.67 3.09 3.43 3.85 3.86 2.89 3.35 3.58 3.68 2.95 3.37 3.69 3.50
CleanMel-S-map (ours) 3.33 3.58 4.11 3.96 3.02 3.36 3.73 3.79 3.39 3.62 4.13 4.08 3.33 3.58 4.10 4.03 3.25 3.55 4.03 3.82
CleanMel-S-mask (ours) 3.26 3.56 4.01 3.81 2.97 3.34 3.68 3.71 3.31 3.58 4.08 4.01 3.25 3.53 4.05 3.97 3.15 3.49 3.93 3.71
CleanMel-L-mask (ours) 3.30 3.59 4.04 3.87 2.99 3.35 3.70 3.74 3.31 3.58 4.08 4.02 3.25 3.53 4.05 4.00 3.15 3.49 3.93 3.71
TABLE V: PESQ scores on test sets with clean reference speech.
Mode Enhancement method CHiME4 simu. DNS w.o. reverb. REVERB simu. RealMAN static
unproc. 1.20 1.58 1.50 1.14
Online FullSubNet [15] 1.80 2.71 2.39 1.36
Demucs [50] 1.64 2.63 1.92 -
oSpatialNet [22] 1.81 2.77 2.56 1.87
CleanMel-S-map (ours) 1.77 2.73 2.63 1.79
CleanMel-S-mask (ours) 1.99 2.82 2.63 1.75
Offline FullSubNet [15] 1.87 2.82 2.48 1.61
VoiceFixer [19] 1.57 1.92 1.67 -
StoRM [51] 1.76 2.74 2.52 1.71
SpatialNet [21] 1.90 2.82 2.87 1.93
SpatialNet-Mamba 1.90 2.90 3.06 2.10
CleanMel-S-map (ours) 2.09 2.95 2.92 2.00
CleanMel-S-mask (ours) 2.17 2.97 2.85 1.95
CleanMel-L-mask (ours) 2.35 3.07 2.97 2.01
@@ V-A1 Comparing the training targets of logMel mapping and Mel ratio mask
For online processing, mapping consistently outperforms mask in DNSMOS, mainly due to the higher residual noise of mask, which can be reflected by the much lower BAK scores of mask on the very noisy CHiME4 and RealMAN test sets. When applying the predicted ratio mask on the noisy spectrogram, if speech is highly contaminated by noise, there will exist certain residual noise even if the predicted error of mask is small. By contrast, mapping directly predicts the logMel-spectrogram of speech, which can avoid such residual noise. However, as shown in Table V, mask achieves higher PESQ scores than mapping on the highly noisy CHiME4 and DNS (w.o. reverb) sets. PESQ measures the perceptual similarity of enhanced speech and reference clean speech. The higher PESQ scores indicates that mask performs better on retrieving the target speech through extracting the target speech from the noisy speech with a mask. Directly mapping may erroneously remove or boost those speech components highly contaminated by noise. The enhanced logMel-spectrograms are transformed to waveforms with a neural vocoder, and the speech quality measured with DNSMOS is more affected by the residual noise caused by mask than the mapping error of logMel-spectrogram. Fig 4 shows an example of the enhanced logMel-spectrogram by mask and mapping, which verifies our discussions.
@@ V-A2 Comparing with oSpatialNet and SpatialNet-Mamba
oSpatialNet and SpatialNet-Mamba adopt the same network architectures and serve as the linear-frequency baselines for the proposed CleanMel networks. For both online and offline processing, the proposed Mel-spectrogram enhancement networks noticeably outperform their linear-frequency baselines in DNSMOS, but not necessarily in PESQ. This phenomenon is consistent with the findings in the VoiceFixer paper [19]. The GAN-based nerual vocoder is good at generating high-quality speech, but will possibly reduce the (PESQ) similarity of the enhance speech with the reference speech.
@@ V-A3 Comparing with online baseline networks
Demucs [50] performs well for denoising but not for dereverberation, as it achieves leading performance in DNSMOS P.835 on the (low-reverberation) CHiME4 and EARS sets, but not for other reverberant test sets. Although it was developed for multi-channel speech enhancement in [22], oSpatialNet also performs very well for single-channel speech enhancement compared with other comparison models. oSpatialNet and the proposed CleanMel network work better especially for dereverberation due to their narrow-band processing of room filter convolution.
@@ V-A4 Comparing with offline baseline networks
VoiceFixer [19] is an important baseline for the proposed CleanMel network, as they both perform Mel-spectrogram enhancement. VoiceFixer adopts an advanced ResUNet. We can see that the proposed network noticeably outperforms VoiceFixer, which verifies the strong capability of the proposed cross-band and narrow-band combination network. StoRM [51] is a diffusion-model-based generative network, which often achieves the SOTA performance for speech enhancement. The proposed network consistently outperforms StoRM. Their performance gaps of DNSMOS scores are not large, but the gaps of PESQ scores are very large. This indicates that although generative model can generate high-quality speech, but its fidelity to the source speech is hard to be ensured. SpatialNet [21], especially its Mamba variant, i.e. SpatialNet-Mamba, as a pure discriminative speech enhancement network, achieves promising performance.
Overall, by combining a powerful discriminative Mel-spectrogram enhancement network and a GAN-based generative neural vocoder, the proposed network achieves new SOTA speech enhancement performance for both online and offline processing, and for both denoising and dereverberation. By scaling up the proposed network, the performance can be further improved. Taking the mask scheme as an example, compared to the small model, the large model noticeably improves the PESQ scores (although not the DNSMOS scores). The performance improvement of large model is clearly audible when listening to the enhanced speech demos.
V-B ASR results
@@
Enhancement method CHiME4 REVERB simu. REVERB real RealMAN static
@@
  • †VoiceFixer is evaluated in the same way as the proposed model, namely the enhanced Mel-spectrogram is directly fed to ASR models. To fit the STFT setups of VoiceFixer, the ASR models used for VoiceFixer were re-trained, which are different from the ones used for other methods. Please refer to Table VII for more information about the re-trained ASR models.
Table VI presents the ASR results. We can observe that the ASR results are consistent with the common finding in the filed [53] that single-channel speech enhancement networks usually do not help for ASR due to the speech artifacts caused by the networks. VoiceFixer performs Mel-spectrogram enhancement using an advanced ResUNet. The poor ASR results of VoiceFixer show that ResUNet brings severe artifacts to the enhanced Mel-spectrogram. StoRM does not work well for ASR, which further verifies that the generated speech has low fidelity to the source speech.
One exception is SpatialNet, especially SpatialNet-Mamba, who show clear positive ASR effects on the REVERB-real and RealMAN sets. We believe that the advantages of SpatialNet lie in its narrow-band processing mechanism: i) narrow-band processing is especially efficient for modeling the much simpler narrow-band room filter convolution; ii) as discussed in [14], narrow-band processing can avoid the so-called wide-band artifacts, such as the blurred and/or wrongly deleted/inserted wide-band spectra, which will be very harmful for ASR.
The proposed CleanMel networks further improve the ASR performance on top of SpatialNet-Mamba. Directly enhancing the Mel-frequency spectra is more optimal for ASR than first enhancing the detailed linear-frequency spectra and then compressing to Mel-frequency. Compared to linear-frequency processing, the lower-dimensional Mel-frequency representation is easier to learn. However, this is more valid for the cross-band processing part, but not for the narrow-band (frequency-wise or dimension-wise) processing part. That is possibly why the proposed CleanMel networks are more advantageous on the denoising task of CHiME4 (more relying on cross-band processing) than the dereverberation task of REVERB (more relying on narrow-band processing). For ASR, Mel ratio mask performs better than logMel mapping, but the performance gaps are not large. This indicates that ASR is less affected by the larger residual noise caused by mask than the prediction error of logMel mapping. Scaling up is very effective for further improving the ASR performance, as shown by the results of the large mask model.
@@
V-C Influence of neural vocoder to ASR performance
TABLE VII: ASR result comparisons between taking as input enhanced Mel-spectrogram and vocoder waveform. WER (%) for the English CHiME4 and REVERB sets, and CER (%) for the Chinese RealMAN set.
Enhancement method CHiME4 REVERB simu. REVERB real RealMAN static

### v2

@@ V Results and Analyses
In this section, we present and analyze speech enhancement performance and ASR results, conduct ablation studies, and compare the model size and computational complexity.
@@ V-A Speech enhancement results
TABLE IV: DNSMOS scores on real and simulated test sets.
Enhancement
method
CHiME4 real EARS
VoiceBank
Demand
REVERB real DNS w. rvb. RealMAN static
P.835 P.808 P.835 P.808 P.835 P.808 P.835 P.808 P.835 P.808 P.835 P.808
OVL SIG BAK OVL SIG BAK OVL SIG BAK OVL SIG BAK OVL SIG BAK OVL SIG BAK
unproc. 1.37 1.99 1.34 2.43 2.02 2.94 2.04 2.80 2.53 3.28 2.87 3.04 1.39 1.74 1.53 2.81 1.42 1.89 1.50 2.74 1.53 2.03 1.75 2.56
Online
FullSubNet [15] 2.41 3.05 3.08 3.22 2.73 3.12 3.55 3.52 2.93 3.30 3.79 3.47 2.72 3.13 3.60 3.61 2.41 2.98 3.15 3.36 2.49 2.97 3.33 3.12
Demucs [2] 2.84 3.12 3.90 3.26 2.94 3.21 3.94 3.38 3.03 3.31 3.97 3.43 2.96 3.25 3.94 3.50 2.50 2.82 3.66 3.17 -
LiSenNet [55] 2.26 2.70 3.31 3.08 2.61 2.93 3.63 3.29 2.87 3.16 3.93 3.42 2.54 3.15 3.31 3.38 2.21 2.56 3.49 3.15 2.32 2.70 3.42 2.96
oSpatialNet [22] 2.61 3.15 3.36 3.52 2.85 3.22 3.65 3.65 2.91 3.26 3.80 3.48 3.03 3.33 3.93 3.84 2.79 3.16 3.70 3.67 2.79 3.19 3.66 3.46
CleanMel-S-map 2.95 3.26 3.86 3.77 2.89 3.24 3.69 3.72 3.02 3.28 4.01 3.65 3.17 3.45 4.02 3.94 3.08 3.36 3.97 3.84 2.90 3.24 3.81 3.52
CleanMel-S-mask 2.74 3.26 3.43 3.59 2.83 3.19 3.64 3.70 2.99 3.31 3.90 3.50 3.07 3.40 3.91 3.80 2.95 3.30 3.81 3.75 2.73 3.18 3.53 3.36
Offline
ConvTasNet [1] 2.79 3.19 3.67 3.12 2.81 3.19 3.62 3.30 3.00 3.31 3.93 3.38 2.95 3.25 3.92 3.29 2.50 2.90 3.51 3.05 2.64 3.06 3.54 2.97
FullSubNet [15] 2.47 3.10 3.11 3.27 2.74 3.16 3.51 3.54 3.03 3.34 3.93 3.52 2.69 3.09 3.60 3.60 2.35 2.91 3.10 3.25 2.47 2.97 3.27 3.13
CMGAN [54] 3.11 3.48 3.82 3.55 3.08 3.38 3.82 3.62 3.16 3.44 4.00 3.44 3.07 3.43 3.81 3.63 3.11 3.47 3.82 3.71 2.93 3.38 3.62 3.34
VoiceFixer [19] 2.95 3.25 3.89 3.65 2.94 3.25 3.83 3.55 3.13 3.43 3.98 3.46 2.88 3.26 3.61 3.62 3.11 3.40 3.97 3.84 -
CDiffuse [56] 2.22 3.22 2.28 2.70 2.50 3.17 2.96 2.78 2.83 3.31 3.60 2.99 1.62 2.03 2.07 2.63 1.98 2.65 2.25 2.83 1.92 2.48 2.59 2.44
StoRM [57] 3.29 3.57 4.03 3.87 2.97 3.37 3.65 3.79 3.20 3.46 4.06 3.61 3.25 3.53 4.01 4.01 3.25 3.53 4.03 3.87 3.04 3.42 3.80 3.68
UNIVERSE++ [58] 3.18 3.43 4.05 3.71 2.97 3.32 3.73 3.63 3.18 3.43 4.07 3.56 3.19 3.50 3.95 3.81 2.74 3.34 3.28 3.54 2.93 3.34 3.68 3.49
SpatialNet [21] 2.76 3.32 3.39 3.51 2.86 3.27 3.58 3.59 3.02 3.33 3.92 3.53 3.07 3.41 3.87 3.84 2.91 3.33 3.68 3.65 2.88 3.32 3.66 3.45
SpatialNet-Mamba 2.94 3.35 3.71 3.65 2.94 3.31 3.69 3.67 3.08 3.36 3.98 3.57 3.09 3.43 3.85 3.86 2.89 3.35 3.58 3.68 2.95 3.37 3.69 3.50
CleanMel-S-map 3.33 3.58 4.11 3.96 3.02 3.36 3.73 3.79 3.26 3.49 4.11 3.63 3.39 3.62 4.13 4.08 3.33 3.58 4.10 4.03 3.25 3.55 4.03 3.82
CleanMel-S-mask 3.26 3.56 4.01 3.81 2.97 3.34 3.68 3.71 3.22 3.47 4.08 3.59 3.31 3.58 4.08 4.01 3.25 3.53 4.05 3.97 3.15 3.49 3.93 3.71
CleanMel-L-map 3.36 3.60 4.13 3.99 3.02 3.37 3.73 3.81 3.26 3.50 4.11 3.63 3.39 3.63 4.14 4.09 3.33 3.57 4.12 4.05 3.23 3.52 4.02 3.82
CleanMel-L-mask 3.30 3.59 4.04 3.87 2.99 3.35 3.70 3.74 3.23 3.48 4.09 3.59 3.31 3.58 4.08 4.02 3.25 3.53 4.05 4.00 3.15 3.49 3.93 3.71
TABLE V: PESQ scores on test sets with clean reference speech.
Enhancement
method
CHiME4
simu.
DNS
w.o. rvb.
VoiceBank
Demand
REVERB
simu.
RealMAN
static
unproc. 1.20 1.58 1.97 1.50 1.14
Online
FullSubNet [15] 1.80 2.71 2.73 2.39 1.54
Demucs [2] 1.64 2.63 2.55 1.92 -
LiSenNet [55] 1.87 2.58 2.72 2.03 1.65
oSpatialNet [22] 1.81 2.77 2.60 2.56 1.87
CleanMel-S-map 1.77 2.73 2.43 2.63 1.79
CleanMel-S-mask 1.99 2.82 2.55 2.63 1.75
Offline
ConvTasNet [1] 1.81 2.56 2.53 2.18 1.55
FullSubNet [15] 1.87 2.82 2.88 2.48 1.61
CMGAN [54] 1.85 2.72 2.53 2.68 1.82
VoiceFixer [19] 1.57 1.92 2.37 1.67 -
CDiffuse [56] 1.54 1.90 2.39 1.42 1.13
StoRM [57] 1.76 2.74 2.73 2.52 1.71
UNIVERSE++ [58] 1.89 2.74 2.87 2.42 1.70
SpatialNet [21] 1.90 2.82 2.71 2.87 1.93
SpatialNet-Mamba 1.90 2.90 2.72 3.06 2.10
CleanMel-S-map 2.09 2.95 2.76 2.92 2.00
CleanMel-S-mask 2.17 2.97 2.81 2.85 1.95
CleanMel-L-map 2.23 3.06 2.85 3.01 2.10
CleanMel-L-mask 2.35 3.07 2.89 2.97 2.01
@@ V-A1 Comparing the training targets of logMel mapping and Mel ratio mask
For online processing, mapping consistently outperforms mask in DNSMOS, mainly due to the higher residual noise of mask, which can be reflected by the much lower BAK scores of mask on the very noisy CHiME4 and RealMAN test sets. When applying the predicted ratio mask on the noisy spectrogram, if speech is highly contaminated by noise, there will exist certain residual noise even if the predicted error of mask is small. By contrast, mapping directly predicts the logMel of speech, which can avoid such residual noise. However, as shown in Table V, mask achieves higher PESQ scores than mapping on the highly noisy CHiME4 and DNS (w.o. reverb) sets. PESQ measures the perceptual similarity of enhanced speech and reference clean speech. The higher PESQ scores indicates that mask performs better on retrieving the target speech through extracting the target speech from the noisy speech with a mask. Directly mapping may erroneously remove or boost those speech components highly contaminated by noise. The enhanced logMel-spectrograms are transformed to waveforms with a neural vocoder, and the speech quality measured with DNSMOS is more affected by the residual noise caused by mask than the mapping error of logMel-spectrogram. Fig. 4 shows an example of the enhanced logMel-spectrogram by mask and mapping, which verifies our discussions.
@@ V-A2 Comparing with oSpatialNet and SpatialNet-Mamba
oSpatialNet and SpatialNet-Mamba adopt the same network architectures and serve as the linear-frequency baselines for the proposed CleanMel networks. For both online and offline processing, the proposed Mel-spectrogram enhancement networks noticeably outperform their linear-frequency baselines in DNSMOS, but not necessarily in PESQ. This phenomenon is consistent with the findings in the VoiceFixer [19]. The GAN-based nerual vocoder is good at generating high-quality speech, but will possibly reduce the similarity of the enhance speech with the reference speech (PESQ).
@@ V-A3 Comparing with online baseline networks
Demucs [2] performs well for denoising but not for dereverberation, as it achieves leading performance in DNSMOS P.835 on the (low-reverberation) CHiME4, EARS and VoiceBank Demand sets, but not for other reverberant test sets. As a light-weight network, LiSenNet [55] achieves promising results. On several test sets, the performance measures of LiSenNet approach even surpass the ones of FullSubNet [15]. Although it was developed for multi-channel speech enhancement in [22], oSpatialNet also performs very well for single-channel speech enhancement compared with other comparison models. oSpatialNet and the proposed CleanMel network work better especially for dereverberation due to their narrow-band processing of room filter convolution.
@@ V-A4 Comparing with offline baseline networks
Among the three discriminative methods (ConvTasNet [1], FullSubNet [15] and CMGAN [54]), the more recently proposed CMGAN performs the best and achieves top-tier performance measures. VoiceFixer [19] is an important baseline for the proposed CleanMel network, as they both perform Mel-spectrogram enhancement. VoiceFixer adopts an advanced ResUNet. We can see that the proposed network noticeably outperforms VoiceFixer, which verifies the strong capability of the proposed cross-band and narrow-band combination network. Among the three diffusion-model-based generative methods (CDiffuse [56], StoRM [57] and UNIVERSE++ [58]), StoRM and UNIVERSE++ perform better than CDiffuse. The proposed network consistently outperforms StoRM. Their performance gaps of DNSMOS scores are not large, but the gaps of PESQ scores are very large. This indicates that although generative model can generate high-quality speech, but its fidelity to the source speech is hard to be ensured. SpatialNet [21], especially its Mamba variant, i.e. SpatialNet-Mamba, as a pure discriminative network, achieves promising results.
Overall, by combining a powerful discriminative Mel-spectrogram enhancement network and a GAN-based generative neural vocoder, the proposed network achieves new SOTA speech enhancement performance for both online and offline processing, and for both denoising and dereverberation. By scaling up the proposed network, the performance can be further improved. Compared to the small models, the large models noticeably improve the PESQ scores (although not the DNSMOS scores).
@@
Enhancement
method
CHiME4 REVERB simu. REVERB real
RealMAN
static
@@
  • †VoiceFixer is evaluated in the same way as the proposed model, namely the enhanced Mel-spectrogram is directly fed to ASR models. To fit the STFT setups of VoiceFixer, the ASR models used for VoiceFixer were re-trained, which are different from the ones used for other methods. Please refer to Table VIII for more information about the re-trained ASR models.
V-B ASR results
Table VI presents the ASR results. We can observe that the ASR results are consistent with the common finding in the field [6] that single-channel speech enhancement networks usually do not help for ASR due to the speech artifacts caused by the networks. VoiceFixer performs Mel-spectrogram enhancement using an advanced ResUNet. The poor ASR results of VoiceFixer show that ResUNet brings severe artifacts to the enhanced Mel-spectrogram. StoRM does not work well for ASR, which further verifies that the generated speech has low fidelity to the source speech.
Three exceptions are CMGAN, SpatialNet and SpatialNet-Mamba. They show clear positive ASR effects on REVERB-real. SpatialNet and SpatialNet-Mamba also show positive ASR effects on RealMAN. We believe that the advantages of CMGAN and SpatialNet lie in their narrow-band processing mechanism (the Conformer-Time block in CMGAN and the narrow-band block in SpatialNet): i) narrow-band processing is especially efficient for modeling the narrow-band room filter convolution; ii) as discussed in [14], narrow-band processing can avoid the so-called wide-band artifacts, such as the blurred and/or wrongly deleted/inserted wide-band spectra, which will be very harmful for ASR.
The proposed CleanMel networks further enhance ASR performance beyond SpatialNet-Mamba. Directly enhancing the Mel-frequency spectra is more optimal for ASR than first enhancing the detailed linear-frequency spectra and then compressing to Mel-frequency. Compared to linear-frequency processing, the lower-dimensional Mel-frequency representation is easier to learn. However, this is more valid for the cross-band processing part, but not for the narrow-band (frequency-wise or dimension-wise) processing part. That is possibly why the proposed CleanMel networks are more advantageous on the denoising task of CHiME4 (relying more on cross-band processing) than the dereverberation task of REVERB (relying more on narrow-band processing). For ASR, Mel ratio mask performs better than logMel mapping, but the performance gaps are not large. This indicates that ASR is less affected by the larger residual noise caused by mask than the prediction error of logMel mapping. Scaling up is very effective for further improving the ASR performance.
@@ V-B ASR results
Figure 5: Comparison of validation losses between linear- and Mel-frequency speech enhancement.
V-C Comparing linear- and Mel-frequency speech enhancement
We have mentioned that, compared to linear-frequency enhancement, Mel-frequency enhancement is easier to learn due to its lower frequency dimension, which is especially valid for the cross-band learning module in the proposed model. To verify this, we compare the validation loss (prediction error) of conducting enhancement in 80, 160 Mel-frequency domain and 256 linear-frequency domain. The CleanMel-S-mask model is adopted. Since the masking ratio (defined in Eq. (4)) has a constant range of values, i.e. [0,1][0,1][ 0 , 1 ], its prediction error in different frequency domains can be directly compared.
The validation losses for the first 100 K training steps are shown in Fig. 5. It can be seen that the prediction errors are indeed lowered as the number of frequencies is decreased. Note that, the computational complexity of enhancement models are roughly proportional to the number of frequencies as well.
V-D Influence of logMel clip value
TABLE VII: The influence of different logMel clip values ($\epsilon$) to speech quality (DNSMOS) and ASR (WER).
Clip values ($\epsilon$)
of clean speech
CHiME4 REVERB
DNSMOS
P.835 OVL
WER
DNSMOS
P.835 OVL
WER
$1e-10$ 3.27 3.11 3.26 3.39
$1e-5$ 3.27 3.16 3.28 3.39
[…]
V-E Influence of neural vocoder to ASR performance
TABLE VIII: ASR result comparisons between taking as input enhanced Mel-spectrogram and vocoder waveform. WER (%) for the English CHiME4 and REVERB sets, and CER (%) for the Chinese RealMAN set.
Enhancement
method
CHiME4 REVERB simu. REVERB real
RealMAN
static